A student performing a physics lab is asked to determine their power output while jogging up a flight of stairs.  If the student’s weight is 820 N and the height of the flight of stairs is 295 cm, how much average power is exerted if the student runs up the stairs in 11 seconds?  How would this relate to the student’s power input (the rate at which stored energy is burned within their body)?


1
Expert's answer
2021-05-10T09:20:21-0400

Find the work required to go 295 cm high:


"W=Fh."

The power describes how quickly the work can be done:


"P=\frac Wt=\frac{Fh}{t}=\frac{820·2.95}{11}=220\text{ W}."

This activity will increase the rate at which stored energy is burned within the student's body.